# Demystify Mamba in Vision: A Linear Attention Perspective

### v1

@@ Abstract
Mamba is an effective state space model with linear computation complexity. It has recently shown impressive efficiency in dealing with high-resolution inputs across various vision tasks. In this paper, we reveal that the powerful Mamba model shares surprising similarities with linear attention Transformer, which typically underperform conventional Transformer in practice. By exploring the similarities and disparities between the effective Mamba and subpar linear attention Transformer, we provide comprehensive analyses to demystify the key factors behind Mamba’s success. Specifically, we reformulate the selective state space model and linear attention within a unified formulation, rephrasing Mamba as a variant of linear attention Transformer with six major distinctions: input gate, forget gate, shortcut, no attention normalization, single-head, and modified block design. For each design, we meticulously analyze its pros and cons, and empirically evaluate its impact on model performance in vision tasks. Interestingly, the results highlight the forget gate and block design as the core contributors to Mamba’s success, while the other four designs are less crucial. Based on these findings, we propose a Mamba-Like Linear Attention (MLLA) model by incorporating the merits of these two key designs into linear attention. The resulting model outperforms various vision Mamba models in both image classification and high-resolution dense prediction tasks, while enjoying parallelizable computation and fast inference speed. Code is available at https://github.com/LeapLabTHU/MLLA.
@@ 1 Introduction
However, Mamba is not the first model to achieve global modeling with linear complexity. Linear attention [20], an early work, was proposed as an computationally efficient alternative to the widely adopted Softmax attention [34], namely dot-product attention. Specifically, linear attention replaces the non-linear Softmax function in attention operation with linear normalization. This enables a change in computation order from $(QK^{\top})V$ to $Q(K^{\top}V)$, thus reducing computation complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. Despite its efficiency, previous works [3, 31, 12] proved that linear attention suffers from insufficient expressive power, making it impractical for real applications. Surprisingly, we find a very close relationship between the formulas of high-performance Mamba and subpar linear attention Transformer. Therefore, a compelling research question emerges: What factors contribute to Mamba’s success and its significant superiority to linear attention Transformer?
In this paper, we offer both theoretical and empirical analyses to unveil Mamba through the lens of linear attention Transformer. Specifically, we rewrite the formulas of selective state space model and linear attention within a unified formulation, depicting Mamba as a variation of linear attention Transformer with six distinctions: input gate, forget gate, shortcut, no attention normalization, single-head, and modified block design. To demystify what factors lead to Mamba’s effectiveness, empirical studies on vision tasks are conducted to assess the impact of each special design. The results demonstrate that the forget gate and block design tend to be the two core contributors to Mamba’s superiority. While the block design can be easily adopted, the forget gate necessitates recurrent computation, which may not be well-suited for non-auto-regressive vision models. Therefore, we delve into the essence of the forget gate and verify that it can be replaced by suitable positional encoding in vision tasks. Based on our findings, we introduce the two core contributors or their alternatives to linear attention Transformer, presenting our Mamba-Like Linear Attention (MLLA) model. Experimental results demonstrate that MLLA achieves superior results to various vision Mamba models in both image classification and high-resolution dense prediction tasks, validating that linear attention can surpass Mamba with the merits of two core designs.
@@
  • We develop a series of linear attention vision Transformer models named MLLA, which inherit the core merits of Mamba and tend to be more suitable for vision tasks than the original Mamba model.
@@ 2 Related Works
Vision Transformer and attention. Originating from natural language processing, Transformer and attention have been highly successful in vision tasks. However, the quadratic complexity of widely adopted Softmax attention [34] poses challenges in handling high-resolution images. Numerous works have attempted to reduce the computational cost by introducing local attention windows [26, 7, 15] or sparsity [35, 37, 47]. Linear attention [20], another approach, inherently offers linear complexity $\mathcal{O}(N)$ and is capable of modeling long sequences. Despite its efficiency, previous works [3, 31, 39, 13] have shown that linear attention always fails to deliver satisfactory results, limiting its applicability.
Mamba [11] is a recently proposed state space model that achieves effective sequence modeling with linear complexity. Motivated by its potential for modeling high-resolution images, many researchers try to apply Mamba to vision tasks [11, 25, 21, 19, 30, 40]. For instance, VMamba [25] introduces a cross-scan module to enable 1D selective scanning in 2D image space. LocalMamba [19] utilizes local windows to enhance local modeling capability. EfficientVMamba [30] designs an atrous-based selective scan approach to enhance efficiency. In addition, MambaOut [42] analyzes whether Mamba is needed for vision, and explainability methods [1] have also been proposed.
@@ 4.1 Interpreting Selective State Space Model as Linear Attention
In addition to these four differences, it is also important to note that eq. 12 represents single-head linear attention. This indicates that the selective state space model is akin to single-head linear attention and does not incorporate a multi-head design.
@@ 4.2 Analysis of Differences in Core Operations
Figure 3: Illustration of the macro designs of linear attention Transformer, Mamba and our MLLA.
@@
Multi-head. Linear attention commonly utilizes multi-head design [34] for better outcome. Multi-head attention allows the model to simultaneously attend to information from various representation subspaces at different positions, thus enhancing its expressive power.
@@ 5.1 Implementation
We employ the widely used Swin Transformer [26] architecture to verify the effects of the six differences. Firstly, we substitute the Softmax attention in Swin Transformer with linear attention to create our baseline model. Subsequently, we separately introduce each distinction to the baseline model to assess its impact. Based on the results, we further validate whether linear attention can achieve superior results with the merits of Mamba. Specifically, we integrate the useful designs into linear attention Transformer to create our Mamba-Like Linear Attention (MLLA) model, and assess its effectiveness by comparing it with various vision Mamba designs across multiple tasks, including ImageNet-1K classification [6], COCO object detection [24], and ADE20K semantic segmentation [46]. Detailed model architectures and training setups are shown in the Appendix.
@@ 5.2 Empirical Analysis of the Differences
Block Design. We employ two ways to assess the effects of Mamba’s block design: 1. Substituting the entire Transformer block with Mamba’s block architecture. 2. Replacing the attention sub-block with Mamba’s block design, while preserving the MLP sub-block. In both settings, the selective SMM in Mamba’s block is substituted with linear attention. To maintain similar FLOPs, we employ Mamba expansion factors [11] $E=2.0$ and $E=1.0$ for the two settings, respectively. The results are presented in Tab. 1 as “Block Design all” and “Block Design sub”. Both replacement approaches result in performance improvements, demonstrating the efficacy of Mamba’s macro design. Substituting the attention sub-block yields better result, which creates our MLLA block shown in Fig. 3(c). Notably, we omit the $V$ projection before linear attention calculation, as a similar input projection already exists. The module complexity of a MLLA block is expressed as:
$\begin{split}\Omega(\text{MLLA})=&\ \underbrace{2NC^{2}+2NC^{2}+NC^{2}}_{\text{\color[rgb]{0,0,1} In/Out,\ Q/K,\ Gate\ Projection}}\ \ +\!\!\!\!\underbrace{2NCd}_{\text{\color[rgb]{0,0,1} Linear\ Attention}}\!\!\!\!+\ \ \underbrace{k^{2}NC}_{\text{\color[rgb]{0,0,1} DWConv}}\ \ +\ \ \underbrace{8NC^{2}}_{\text{\color[rgb]{0,0,1} MLP}}\ ,\end{split}$ (13)
@@ 5.3 Comparison with Mamba in Vision
Based on our findings, we integrate the forget gate and block design into linear attention, introducing our MLLA model. Notably, we practically use LePE, CPE, and RoPE to replace the forget gate’s local bias, input-dependent positional information, and global positional information, respectively.
ImageNet classification. As shown in Tab. 3, our MLLA models consistently outperform various vision Mamba models across all model sizes, owing to the integration of useful designs from both Mamba and linear attention. These results also validate that with the merits of Mamba’s two key designs, the inferior linear attention Transformer can surpass high-performance Mamba. Notably, we empirically observe that MLLA exhibits greater scalability compared to vision Mamba models, as MLLA-B achieves an accuracy of 85.3, surpassing other models by a significant margin. Additionally, MLLA also outperforms various CNN and vision Transformer designs. For instance, MLLA exhibits better performance than MambaOut [42], a recent work that removes the selective SSM in Mamba and employs a gated convolution architecture.
@@
Inference time. We offer real speed measurements in Fig. 6. Substituting the forget gate with positional encoding, our MLLA models benefit from parallelizable computation, resulting in significantly faster inference speeds compared to vision Mamba models. For instance, our model achieves 4.5x faster inference speed than Mamba2D [21], while maintaining better accuracy. Compared to the highly optimized VMamba model [25], our model also delivers a 1.5x speedup accompanied by a 0.5 accuracy gain. These substantial improvements in model speed further support our analysis that the parallelizable MLLA is more suitable than Mamba for modeling non-causal data such as images.
@@
MLLA-B 128M 1183G 51.9 52.5
COCO object detection. As shown in Tab. 4, on the COCO dataset, MLLA models also achieve superior results to vision Mamba models, implying their effectiveness in high-resolution dense prediction tasks. MLLA offers effective global modeling with linear complexity $\mathcal{O}(N)$ (see eq. 13) and parallelizable computation, making it ideally suitable for high-resolution image modeling scenarios. Notably, MLLA outperforms MambaOut [42] by a significant margin, which aligns with the findings in MambaOut [42].
ADE-20K semantic segmentation. We report the results on ADE-20K dataset in Tab. 5, where “SS” and “MS” denote single-scale and multi-scale testing, respectively. Similar to the object detection task, MLLA yields better results in semantic segmentation, further verifying our analyses and the effectiveness of MLLA model.
@@ 6 Conclusion
This paper reveals the surprisingly close relationship between the powerful Mamba and subpar linear attention Transformer, shedding some light on Mamba’s superiority and success. We rephrase Mamba as a variant of linear attention Transformer and identify its six major special designs: input gate, forget gate, shortcut, no attention normalization, single-head and modified block design. Empirical validation shows that the forget gate and block design largely enhance performance, while the other distinctions offer marginal contributions or impair model performance. Based on our findings, we propose our Mamba-Like Linear Attention (MLLA) model by incorporating the merits of these two key designs into linear attention. MLLA surpasses various vision Mamba models across multiple tasks, while maintaining parallel computation and high inference speed.
@@ Appendix A Datasets and Experiment Details
ImageNet classification. The ImageNet-1K dataset comprises 1.28 million training images and 50,000 validation images, encompassing 1,000 classes. For a fair comparison, we train our models under the same settings as Swin Transformer [26]. Specifically, we utilize AdamW [28] optimizer to train all our models from scratch for 300 epochs. We apply a cosine learning rate decay schedule with a linear warm-up of 20 epochs and a weight decay of 0.05. The total batch size is 4096 and initial learning rate is set to $4\times{10}^{-3}$. Augmentation and regularization strategies includes RandAugment [5], Mixup [44], CutMix [43], and random erasing [45]. In the training of MLLA models, MESA [9] is employed to prevent overfitting.
COCO object detection. COCO [24] dataset is a widely adopted benchmark for object detection and instance segmentation with 118K training and 5K validation images. We follow the standard 1x and 3x Mask R-CNN [16] training setting in Swin Transformer [26] to conduct our experiments. The pretrained MLLA models are employed as backbones.
@@ Appendix B Additional Experimental Results
Additional comparison with advanced linear attention designs. The results are shown in Tab. 6. We empirically find that MLLA outperforms various advanced linear attention designs without bells and whistles.
@@
MLLA (Ours) 25M 4.2G 83.5
@@ Appendix C Model Architectures
We illustrate the architecture of our MLLA model in Fig. 7 and summarize the detailed structure in Tab. 7. We adopt the common 4-stage framework to build MLLA model by stacking our MLLA blocks at each stage.
@@ Appendix D Limitations
In this paper, we explore the similarities and disparities between Mamba and linear attention Transformer, providing comprehensive analyses to demystify the key factors behind Mamba’s success. Specifically, we begin with the formulas and rephrase Mamba as a variant of linear attention Transformer with six major distinctions: input gate, forget gate, shortcut, no attention normalization, single-head and modified block design. Moreover, we meticulously analyze the pros and cons of each design and prove that the forget gate and block design are the core contributors to Mamba’s success. Based on our findings, we propose our Mamba-Like Linear Attention (MLLA) model, which surpasses various vision Mamba models across multiple tasks, while maintaining parallel computation and high inference speed. However, there may be other small differences between the implementation details of Mamba and linear attention Transformer, and this paper is not exhaustive.
Figure 7: The architecture of MLLA model.
Table 7: Architectures of MLLA models.
stage output MLLA-T MLLA-S MLLA-B

### v2

@@ Abstract
Mamba is an effective state space model with linear computation complexity. It has recently shown impressive efficiency in dealing with high-resolution inputs across various vision tasks. In this paper, we reveal that the powerful Mamba model shares surprising similarities with linear attention Transformer, which typically underperform conventional Transformer in practice. By exploring the similarities and disparities between the effective Mamba and subpar linear attention Transformer, we provide comprehensive analyses to demystify the key factors behind Mamba’s success. Specifically, we reformulate the selective state space model and linear attention within a unified formulation, rephrasing Mamba as a variant of linear attention Transformer with six major distinctions: input gate, forget gate, shortcut, no attention normalization, single-head, and modified block design. For each design, we meticulously analyze its pros and cons, and empirically evaluate its impact on model performance in vision tasks. Interestingly, the results highlight the forget gate and block design as the core contributors to Mamba’s success, while the other four designs are less crucial. Based on these findings, we propose a Mamba-Inspired Linear Attention (MILA) model by incorporating the merits of these two key designs into linear attention. The resulting model outperforms various vision Mamba models in both image classification and high-resolution dense prediction tasks, while enjoying parallelizable computation and fast inference speed. Code is available at https://github.com/LeapLabTHU/MLLA.
@@ 1 Introduction
However, Mamba is not the first model to achieve global modeling with linear complexity. Linear attention [26], an early work, was proposed as an computationally efficient alternative to the widely adopted Softmax attention [42], namely dot-product attention. Specifically, linear attention replaces the non-linear Softmax function in attention operation with linear normalization. This enables a change in computation order from $(QK^{\top})V$ to $Q(K^{\top}V)$, thus reducing computation complexity from $\mathcal{O}(N^{2})$ to $\mathcal{O}(N)$. Despite its efficiency, previous works [4, 39, 15, 16] proved that linear attention suffers from insufficient expressive power, making it impractical for real applications. Surprisingly, we find a very close relationship between the formulas of high-performance Mamba and subpar linear attention Transformer. Therefore, a compelling research question emerges: What factors contribute to Mamba’s success and its significant superiority to linear attention Transformer?
In this paper, we offer both theoretical and empirical analyses to unveil Mamba through the lens of linear attention Transformer. Specifically, we rewrite the formulas of selective state space model and linear attention within a unified formulation, depicting Mamba as a variation of linear attention Transformer with six distinctions: input gate, forget gate, shortcut, no attention normalization, single-head, and modified block design. To demystify what factors lead to Mamba’s effectiveness, empirical studies on vision tasks are conducted to assess the impact of each special design. The results demonstrate that the forget gate and block design tend to be the two core contributors to Mamba’s superiority. While the block design can be easily adopted, the forget gate necessitates recurrent computation, which may not be well-suited for non-auto-regressive vision models. Therefore, we delve into the essence of the forget gate and verify that it can be replaced by suitable positional encoding in vision tasks. Based on our findings, we introduce the two core contributors or their alternatives to linear attention Transformer, presenting our Mamba-Inspired Linear Attention (MILA) model. Experimental results demonstrate that MILA achieves superior results to various vision Mamba models in both image classification and high-resolution dense prediction tasks, validating that linear attention can surpass Mamba with the merits of two core designs.
@@
  • We develop a series of linear attention vision Transformer models named MILA, which inherit the core merits of Mamba and tend to be more suitable for vision tasks than the original Mamba model.
@@ 2 Related Works
Vision Transformer and attention. Originating from natural language processing, Transformer and attention have been highly successful in vision tasks [32, 3, 36, 37], demonstrating superiority to the conventional CNN models[18, 7, 12]. However, the quadratic complexity of widely adopted Softmax attention [42] poses challenges in handling high-resolution images. Numerous works have attempted to reduce the computational cost by introducing local attention windows [32, 9, 20] or sparsity [43, 45, 56]. Linear attention [26], another approach, inherently offers linear complexity $\mathcal{O}(N)$ and is capable of modeling long sequences. Despite its efficiency, previous works [4, 39, 47, 17] have shown that linear attention always fails to deliver satisfactory results, limiting its applicability.
Mamba [14] is a recently proposed state space model that achieves effective sequence modeling with linear complexity. Motivated by its potential for modeling high-resolution images, many researchers try to apply Mamba to vision tasks [14, 31, 27, 25, 38, 49, 48, 23]. For instance, VMamba [31] introduces a cross-scan module to enable 1D selective scanning in 2D image space. LocalMamba [25] utilizes local windows to enhance local modeling capability. EfficientVMamba [38] designs an atrous-based selective scan approach to enhance efficiency. In addition, MambaOut [51] analyzes whether Mamba is needed for vision, and explainability methods [1] have also been proposed.
@@ 4.1 Interpreting Selective State Space Model as Linear Attention
In addition to these four differences, it is also important to note that eq. 12 represents single-head linear attention as there is only one group of $Q,K$. This indicates that the selective state space model is akin to single-head linear attention and does not incorporate a multi-head design.
@@ 4.2 Analysis of Differences in Core Operations
Figure 3: Illustration of the macro designs of linear attention Transformer, Mamba and our MILA.
@@
Multi-head. Linear attention commonly utilizes multi-head design [42] for better outcome. Multi-head attention employs multiple groups of $Q,K$ to produce attention matrices and allows the model to simultaneously attend to information from various representation subspaces at different positions, thus enhancing its expressive power.
@@ 5.1 Implementation
We employ the widely used Swin Transformer [32] architecture to verify the effects of the six differences. Firstly, we substitute the Softmax attention in Swin Transformer with linear attention to create our baseline model. Subsequently, we separately introduce each distinction to the baseline model to assess its impact. Based on the results, we further validate whether linear attention can achieve superior results with the merits of Mamba. Specifically, we integrate the useful designs into linear attention Transformer to create our Mamba-Inspired Linear Attention (MILA) model, and assess its effectiveness by comparing it with various vision Mamba designs across multiple tasks, including ImageNet-1K classification [8], COCO object detection [30], and ADE20K semantic segmentation [55]. Detailed model architectures and training setups are shown in the Appendix.
@@ 5.2 Empirical Analysis of the Differences
Block Design. We employ two ways to assess the effects of Mamba’s block design: 1. Substituting the entire Transformer block with Mamba’s block architecture. 2. Replacing the attention sub-block with Mamba’s block design, while preserving the MLP sub-block. In both settings, the selective SMM in Mamba’s block is substituted with linear attention. To maintain similar FLOPs, we employ Mamba expansion factors [14] $E=2.0$ and $E=1.0$ for the two settings, respectively. The results are presented in Tab. 1 as “Block Design all” and “Block Design sub”. Both replacement approaches result in performance improvements, demonstrating the efficacy of Mamba’s macro design. Substituting the attention sub-block yields better result, which creates our MILA block shown in Fig. 3(c). Notably, we omit the $V$ projection before linear attention calculation, as a similar input projection already exists. The module complexity of a MILA block is expressed as:
$\begin{split}\Omega(\text{MILA})=&\ \underbrace{2NC^{2}+2NC^{2}+NC^{2}}_{\text{\color[rgb]{0,0,1} In/Out,\ Q/K,\ Gate\ Projection}}\ \ +\!\!\!\!\underbrace{2NCd}_{\text{\color[rgb]{0,0,1} Linear\ Attention}}\!\!\!\!+\ \ \underbrace{k^{2}NC}_{\text{\color[rgb]{0,0,1} DWConv}}\ \ +\ \ \underbrace{8NC^{2}}_{\text{\color[rgb]{0,0,1} MLP}}\ ,\end{split}$ (13)
@@ 5.3 Comparison with Mamba in Vision
Based on our findings, we integrate the forget gate and block design into linear attention, introducing our MILA model. Notably, we practically use LePE, CPE, and RoPE to replace the forget gate’s local bias, input-dependent positional information, and global positional information, respectively.
ImageNet classification. As shown in Tab. 3, our MILA models consistently outperform various vision Mamba models across all model sizes, owing to the integration of useful designs from both Mamba and linear attention. These results also validate that with the merits of Mamba’s two key designs, the inferior linear attention Transformer can surpass high-performance Mamba. Notably, we empirically observe that MILA exhibits greater scalability compared to vision Mamba models, as MILA-B achieves an accuracy of 85.3, surpassing other models by a significant margin. Additionally, MILA also outperforms various CNN and vision Transformer designs. For instance, MILA exhibits better performance than MambaOut [51], a recent work that removes the selective SSM in Mamba and employs a gated convolution architecture.
@@
Inference time. We offer real speed measurements in Fig. 6. Substituting the forget gate with positional encoding, our MILA models benefit from parallelizable computation, resulting in significantly faster inference speeds compared to vision Mamba models. For instance, our model achieves 4.5x faster inference speed than Mamba2D [27], while maintaining better accuracy. Compared to the highly optimized VMamba model [31], our model also delivers a 1.5x speedup accompanied by a 0.5 accuracy gain. These substantial improvements in model speed further support our analysis that the parallelizable MILA is more suitable than Mamba for modeling non-causal data such as images.
@@
MILA-B 128M 1183G 51.9 52.5
COCO object detection. As shown in Tab. 4, on the COCO dataset, MILA models also achieve superior results to vision Mamba models, implying their effectiveness in high-resolution dense prediction tasks. MILA offers effective global modeling with linear complexity $\mathcal{O}(N)$ (see eq. 13) and parallelizable computation, making it ideally suitable for high-resolution image modeling scenarios. Notably, MILA outperforms MambaOut [51] by a significant margin, which aligns with the findings in MambaOut [51].
ADE-20K semantic segmentation. We report the results on ADE-20K dataset in Tab. 5, where “SS” and “MS” denote single-scale and multi-scale testing, respectively. Similar to the object detection task, MILA yields better results in semantic segmentation, further verifying our analyses and the effectiveness of MILA model.
@@ 6 Conclusion
This paper reveals the surprisingly close relationship between the powerful Mamba and subpar linear attention Transformer, shedding some light on Mamba’s superiority and success. We rephrase Mamba as a variant of linear attention Transformer and identify its six major special designs: input gate, forget gate, shortcut, no attention normalization, single-head and modified block design. Empirical validation shows that the forget gate and block design largely enhance performance, while the other distinctions offer marginal contributions or impair model performance. Based on our findings, we propose our Mamba-Inspired Linear Attention (MILA) model by incorporating the merits of these two key designs into linear attention. MILA surpasses various vision Mamba models across multiple tasks, while maintaining parallel computation and high inference speed.
Acknowledgement
This work is supported in part by the National Natural Science Foundation of China under Grants 42327901 and 62321005.
@@ Appendix A Datasets and Experiment Details
ImageNet classification. The ImageNet-1K dataset comprises 1.28 million training images and 50,000 validation images, encompassing 1,000 classes. For a fair comparison, we train our models under the same settings as Swin Transformer [32]. Specifically, we utilize AdamW [34] optimizer to train all our models from scratch for 300 epochs. We apply a cosine learning rate decay schedule with a linear warm-up of 20 epochs and a weight decay of 0.05. The total batch size is 4096 and initial learning rate is set to $4\times{10}^{-3}$. Augmentation and regularization strategies includes RandAugment [6], Mixup [53], CutMix [52], and random erasing [54]. In the training of MILA models, MESA [11] is employed to prevent overfitting.
COCO object detection. COCO [30] dataset is a widely adopted benchmark for object detection and instance segmentation with 118K training and 5K validation images. We follow the standard 1x and 3x Mask R-CNN [21] training setting in Swin Transformer [32] to conduct our experiments. The pretrained MILA models are employed as backbones.
@@ Appendix B Additional Experimental Results
Additional comparison with advanced linear attention designs. The results are shown in Tab. 6. We empirically find that MILA outperforms various advanced linear attention designs without bells and whistles.
@@
MILA (Ours) 25M 4.2G 83.5
Ablation on the impact of MESA. Like in the early stages of studies on visual Transformer, currently vision Mamba research does not have a well-established and universally accepted training protocol. The conventional training setting for vision Transformer may not be optimal for vision Mamba and our Mamba-Inspired Linear Attention. Therefore, in the training of MILA models, we additionally employ the overfitting prevention strategy MESA [11] to fully demonstrate its potential. In Tab. 7, we provide the results without MESA. We can observe that: (1) MESA provides a modest accuracy gain of 0.1-0.3. (2) Without MESA, MILA models still significantly surpass various SOTA vision Mamba models.
Table 7: MILA models trained without MESA.
Model #Params FLOPs Acc.
Vim-S [57] 26M 5.1G 80.3
VMamba-T [31] 31M 4.9G 82.5
LocalVMamba-T [25] 26M 5.7G 82.7
MILA-T (w/o MESA) 25M 4.2G 83.3
MILA-T (w/ MESA) 25M 4.2G 83.5
VMamba-S [31] 50M 8.7G 83.6
LocalVMamba-S [25] 50M 11.4G 83.7
MILA-S (w/o MESA) 43M 7.3G 84.2
MILA-S (w/ MESA) 43M 7.3G 84.3
Mamba2D-B [27] 94M - 82.3
VMamba-B [31] 89M 15.4G 83.9
MILA-B (w/o MESA) 96M 16.2G 85.0
MILA-B (w/ MESA) 96M 16.2G 85.3
@@ Appendix C Model Architectures
We illustrate the architecture of our MILA model in Fig. 7 and summarize the detailed structure in Tab. 8. We adopt the common 4-stage framework to build MILA model by stacking our MILA blocks at each stage.
@@ Appendix D Limitations
In this paper, we explore the similarities and disparities between Mamba and linear attention Transformer, providing comprehensive analyses to demystify the key factors behind Mamba’s success. Specifically, we begin with the formulas and rephrase Mamba as a variant of linear attention Transformer with six major distinctions: input gate, forget gate, shortcut, no attention normalization, single-head and modified block design. Moreover, we meticulously analyze the pros and cons of each design and prove that the forget gate and block design are the core contributors to Mamba’s success. Based on our findings, we propose our Mamba-Inspired Linear Attention (MILA) model, which surpasses various vision Mamba models across multiple tasks, while maintaining parallel computation and high inference speed. However, there may be other small differences between the implementation details of Mamba and linear attention Transformer, and this paper is not exhaustive.
Figure 7: The architecture of MILA model.
Table 8: Architectures of MILA models.
stage output MILA-T MILA-S MILA-B